OpenStudy (kamille):

i need to solve for x: \[x+\lg(1+2^{x})=xlg5+\lg6\]

OpenStudy (kamille):

@amistre64

OpenStudy (anonymous):

x+g6l2x+2xg6l2x+g6l

OpenStudy (kamille):

maybe: \[x+\lg(1+2^{x})=\lg5^{x}+\lg6\]

OpenStudy (kamille):

then: \[x+\lg(1+2^{x})=\lg(5^{x}*6)\] but what to do next? @phi?

OpenStudy (phi):

what base is lg ?

OpenStudy (kamille):

\[x=\lg(\frac{ 6*5^{x} }{1+2^{x} })\]

OpenStudy (kamille):

lg is \[\log _{10}\]

OpenStudy (kamille):

@beneemi I am sorry, can you explain what you have done?

OpenStudy (anonymous):

a friend did this answer i have know clue sorry

OpenStudy (kamille):

oh, it doesnt make sense to me:/

OpenStudy (phi):

you could undo the log 10^x = 6 * 5^x / (1+2^x) and write 10^x as (2*5)^x = 2^x * 5^x

OpenStudy (kamille):

thank you, @phi

OpenStudy (phi):

I get \[ 2^x \cdot 5^x = \frac{6 \cdot 5^x}{(1+2^x)} \] divide both sides by 5^x \[ 2^x = \frac{6 }{(1+2^x)} \] multiply through by 1+2^x \[ 2^{2x} +2^x -6 =0 \] or \[ \left(2^x\right)^2 + 2^x - 6 =0 \] this is a quadratic with roots -3 and 2

OpenStudy (phi):

we can discard the -3. we are left with 2^x = 2 and x= 1